A wire of length 0.50 m is fixed horizontally between two supports separated by 0.50
m. When a mass of 8.0 kg hangs from the middle of the wire, the mid-point sags by
1.00 cm. The diameter of the wire is 2.8 mm. Calculate the Young’s modulus of the
wire. (Given g = 9.81 m s2
)
1
Expert's answer
2021-08-26T19:04:08-0400
"E=\frac{4mgl}{e\pi d^2}\\E=\frac{4(8)(9.81)(0.5)}{(0.01)\pi (0.0028)^2}=640\ MPa"